Reports 1-1 of 1 Clear search Modify search
MIF (ASC)
hirose.chiaki - 1:53 Wednesday 19 April 2023 (24855) Print this report
Increased the gain of YAW in ETM WFS, and succeeded in closing the YAW control loop

Ushiba, K.Tanaka, Hirose

Continued from klog24859.
We tried YAW of WFS, set the filters, actually turned on the controls, adjusted the controls, and found that the YAW of WFS could go in.
We need to adjust the PIT and YAW filters (especially the peaks that appear in the error signal) a bit more.

What we did.

  • Error signal of alignment by ETMX,ETMY in TypeA suspension has peaks at 0.5Hz, 1.69Hz, and 3.1Hz for a 3-stage mode. We set COMP filter and notch filter targeting UGF 1Hz. The model refers to MI_OPLEV's ETMX.
  • Turned on the control and temporarily put in the filter gain adjustment and notch filter during the timescale observation and power spectrum measurement. YAW control was turned on and 4 feedbacks(including PIT) were turned on at the same time.
  • Fig1/This power spectrum is the result of an experiment in progress when WFS ON/OFF. The final results were not measured well, so I will re-measure and post them later.
    /users/Commissioning/data/ASC/2023/0418/SPEC_ETM_WFS_ASCONOFF.xml 
    The filters I put in are as follows.(Fig2/The filter of DSOFT and CSOFT in MEDM screen)
    DoF Added Frequency of Notch filter[Hz]
    DSOFT_PIT 1.44, 1.75, 7.46, 7.398
    CSOFT_PIT 1.44, 1.75, 7.46, 7.398
    DSOFT_YAW 1.44, 2.23, 4.07, 3,14,1.68
    CSOFT_YAW 1.44, 2.23, 4.07, 3,14,1.68
  • Added the following statuses and functions to ASC_LOCK guardian.(Fig3/ The graph of ASC_LOCK guardian)
    [ETM_WFS_ON_WITHOUT_DC (1022) status] This is for turning on WFS without DC.
    • Turn off the filtering of DC in WFS.
    • Turn on 1/2 _times filter in MI_OPLEV. (The contents of MI_OPLEV and WFS filters are different, but once locked, the two controls interfered with each other and the lock became unstable, so 1/2 filter is added to MI_OPLEV. Also, since U.G.F. of  EX and EY in MI_OPLEV control are the same, they are both placed 1/2).
    • Turn off the null of DSOFT and CSOFT_PITYAW. (Turn on the control).
  • [ETM_WFS_DC_ON (1024) status] This is for turning on WFS DC.
    • Put null in MI_OPLEV's filter (turn it off.)
    • Turn ON the filter of DC.
    • Turn OFF 1/2 filter to MI_OPLEV.
  • [DOWN status]
    • Turns on the DSOFT and CSOFT null.(Turn off the control)
    • Set DSOFT and CSOFT gain to 0 to set the OUTPUT signal to 0.
    • After setting the OUTPUT signal to 0, do a clear history and set gain1.
    • Turn off the null filter of MI_OPLEV.(Turn on the control)

Future Plans

  • Instead of open-loop transfer function measurements, excite at a single frequency and measure the phase at that frequency. By comparing the phase of this measurement with the phase of the model, we will estimate the UGF and adjust the gain.
  • Adjustment of notch filters, etc.
  • Investigation of PITYAW coupling of the optical axis, etc.
Images attached to this report
Comments to this report:
hirose.chiaki - 11:07 Thursday 20 April 2023 (24896) Print this report

I'm sorry, The name "MI_OPLEV" is wrong. The correct name is "MN_OPLEV". This filter is to feedback from OPLEV to marionette(MN) stage.
 

Search Help
×

Warning

×